OpenStudy (anonymous):

PRE-CALC/TRIG: Eliminate the parameter and identify the curve: x-2t-3, y=9-4t, 3

OpenStudy (loser66):

tell me what you get after eliminate the t?

OpenStudy (anonymous):

I am trying to help my daughter. she is having trouble eliminating it with the inequalities Can you explain how to do that please

OpenStudy (loser66):

ok, isolate the t from x and y equations, and then, let them = ( because t = .... something to x; and t = ... something to y, so that t is equal, it means something x = something y) then isolate y = another thing to x only, so you have the curve.

OpenStudy (anonymous):

Hold on a sec, I am going to give her my computer and then I won't confuse her....

OpenStudy (loser66):

you have x = 2t -3 ---> 2t = x+3 y = 9-4t ---> 2t = (9 - y)/2 ------------------------------------ x+3 = (9 -y)/2 are you with me then?

OpenStudy (anonymous):

I am having trouble with the inequality part

OpenStudy (loser66):

so , you get the curve?

OpenStudy (loser66):

tell me what you get? the inequality part is the condition (or boundary of the curve/ limits of the curve) solve it later

OpenStudy (loser66):

am I talking with the student or the parent of him/ her?

OpenStudy (anonymous):

student nsorry

OpenStudy (loser66):

ok, write out the curve you get.

OpenStudy (anonymous):

y=9-4(x+3/2)

OpenStudy (loser66):

have to say "wrong", redo

OpenStudy (anonymous):

I do not understand what I did wrong =(

OpenStudy (loser66):

\[x+3= \frac{9-y}{2}\]---> \[2(x+3) = 9-y\] --> 2x +6 = 9-y ---> 2x +6-9=-y ---> -y = 2x-3 ---> y = -2x+3

OpenStudy (loser66):

got what I mean?

OpenStudy (anonymous):

Yes that is so much clearer. thank you!!!

OpenStudy (loser66):

now, the unequality parts. you have x= 2t -3. if t =3 you have x = 3, if t =5 , x = 7 so to the new curve you have, the condition must follow the new variable, that x. so you have the new limits is 3<=x<=7 got it?

OpenStudy (anonymous):

So i just replace the t in the inequality with the new t to find the new limits?

OpenStudy (loser66):

sure. just x, not y because y respect to x, it naturally follow x changing.

OpenStudy (anonymous):

Okay thank you so much for your help!

OpenStudy (loser66):

If you have something which you cannot get from my English, ask your mom/dad. I think he/she can explain you. hehe.. sorry for my bad English.

OpenStudy (anonymous):

Thanks have a blessed day

OpenStudy (loser66):

ty